A ray of Light at an angle of 60 degree on one face of a prism angle 30 degree. The ray emerging out of the prism makes an angle of 30 degree with the incident ray. Calculate the refractive index of the material of the prism.
1
Expert's answer
2013-01-30T08:41:33-0500
On the entrance to the prism,
sin (60 deg) = n * sin (alpha)
On the exit, n*sin(beta) = sin (gamma)

from knowing the prism angle,
60 deg = 30 deg + gamma
from where,
gamma = 30 deg
beta = 60 - alpha.

The resulting system is
sin (60 deg) = n * sin (alpha)
sin (30 deg) = n * sin (60 deg - alpha)

From solving it,
n = sqrt((4+sqrt(3))/3) = 1.3823...

Answer: 1.38.